# Improving Group Robustness on Spurious Correlation Requires Preciser Group Inference

### v1

@@ C.5 Group Distribution Discrepancy Analysis
Figure 13: The result after adjusting the group distribution using the group labels predicted by GIC.
@@
C.6 GIC compute resources and training time
@@
Regarding the runtime of GIC, additional computational overhead compared to methods inferring spurious labels via ERM stems from inputting data representations $\mathbf{z}$ into the GIC model $f_{\mathrm{GIC}}$ for both training and predicting the spurious feature labels $\hat{y}^{tr}_{s}$. This supplementary step should not necessitate substantial computation cost. As the dimensionality of data representation $\mathbf{z}$ is reduced relative to the raw data, and the experiments have evidenced that a simple single-layer neural network for $f_{\mathrm{GIC}}$ suffices in delivering promising performance. Concurrently, it’s noted in Section B.2.3 that the training epoch $K$ can be contained within 20 epochs for all four experimental datasets. We further calculate the total time required for training the spurious feature classifier $f_{\mathrm{GIC}}$ and predicting the spurious feature labels $\hat{y}^{tr}_{s}$ across the four experimental datasets, manifesting that the computational cost introduced by GIC is rather minimal.
Table 6: The average runtime for training $f_{\mathrm{GIC}}$ and predicting $\hat{y}^{tr}_{s}$ over three seeds. The timing commences subsequent to obtaining the data representation $\mathbf{z}$ and concludes upon acquiring the predicted $\hat{y}^{tr}_{s}$.

### v2

@@ C.6 Group Distribution Discrepancy Analysis
Table 7: The result after adjusting the group distribution using the group labels predicted by GIC.
@@
C.7 Compute Resources and Training Time of GIC
@@
Regarding the runtime of GIC, additional computational overhead compared to methods inferring spurious labels via ERM stems from inputting data representations $\mathbf{z}$ into the GIC model $f_{\mathrm{GIC}}$ for both training and predicting the spurious feature labels $\hat{y}^{c}_{s,\mathbf{w}}$. This supplementary step should not necessitate substantial computation cost. As the dimensionality of data representation $\mathbf{z}$ is reduced relative to the raw data, and the experiments have evidenced that a simple single-layer neural network for $f_{\mathrm{GIC}}$ suffices in delivering promising performance. Concurrently, it’s noted in Section B.2.3 that the training epoch $K$ can be contained within 20 epochs for all four experimental datasets. We further calculate the total time required for training the spurious feature classifier $f_{\mathrm{GIC}}$ and predicting the spurious feature labels $\hat{y}^{c}_{s,\mathbf{w}}$ across the four experimental datasets, manifesting that the computational cost introduced by GIC is rather minimal.
Table 8: The average runtime for training $f_{\mathrm{GIC}}$ and predicting $\hat{y}^{c}_{s,\mathbf{w}}$. The timing commences subsequent to obtaining the data representation $\mathbf{z}$ and concludes upon acquiring the predicted $\hat{y}^{c}_{s,\mathbf{w}}$.